OpenStudy (anonymous):

Write an equation of the parabola that opens upward, has a vertex at the orgin , and a focus at (0,1)

OpenStudy (anonymous):

Remember, vertex at the origin means: x^2 = 4py Focus for a vertical parabola is: (h,k+p) = (0,1) Since k is 0, solve for p: 0 + p = 1 p = 1 The equation is therefore: x^2 = 4y x^2/4 = y (C)

OpenStudy (anonymous):

x^2/4=3.16?

OpenStudy (anonymous):

@student365 what are you answer choices

OpenStudy (anonymous):

A. y=1/4x^2 B. y =1/12x^2 c. y= -1/4x^2 d. y=-1/12x^2

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw